# Solving a Stackelberg Game on Transportation Networks in a Dynamic Crime Scenario: A Mixed Approach on Multi-Layer Networks

### v2

@@ 1 Introduction
We consider a Stackelberg game where the defender is the leader, and the attacker is the follower. The assumption is that the defenders have information only about the crime location in a large transportation network. The defenders attempt to capture the attacker before he can flee the city. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model. For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location. For the attacker, the initial strategy is any random path from the crime location to any exit point. An exact algorithm is developed to generate the optimal path for the attacker using Dijkstra’s algorithm on the time-expanded network. The new defender strategy is generated using a greedy algorithm on the time-expanded network. The defender’s mixed strategy set is updated based on the recalculated utility. The defender then selects the best strategy to commit to. This best defender strategy is given as input to the attacker. The attacker tries to generate a new optimal strategy using the exact approach, taking the new defender strategy as input. The attacker updates the pure attacker strategy set, where all strategies are of equal probability. The algorithm checks if, for a certain number of consecutive iterations, the game utility is below a threshold value. If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime. To set a benchmark, the developed Stackelberg game is solved using the adopted MILP approaches for both the attacker and defender.
@@
Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task. We can summarize the main contributions of this paper as follows: (1) To the best of our knowledge, this is the first attempt to solve a Stackelberg game considering the escape interdiction game (EIG) problem using the multi-layer network concept, with a focus on developing solution methodologies. (2) We develop an optimal approach for the attacker and a near-optimal approach for the defenders, and we are not aware of any previous implementations for escape interdiction problems on multi-layer network. (3) Using grid datasets, we demonstrate that our method significantly outperforms the MILP approach in both computational time and solution quality, with minimal performance trade-offs.
@@ 2 Related work
Security games play an important role in providing social security ([23], [25]). Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society. For example, [4] consider search games (SEG) on directed graphs. They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location. They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets. Similarly, [9] consider a leader-follower game and formulate the problem as a mathematical programming model. They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower. Again, [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players. Similarly, [10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution. In the same vein, [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes.
@@
3 Problem description
We consider a two-player Stackelberg game where this sequential game is played between multiple defenders $(D)=\{d_{r}\mid r\in R\}$ and one attacker $(A)$. The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$. $S$ and $A$ represent the sets of probability distributions over the defender and attacker actions, respectively. The game is modeled as a general-sum game, where $x\in S$ and $y\in A$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively. The game begins at time 0 and ends at time $t_{max}>0$. $v_{\infty}$ signifies any exit node in the considered network.
The sequence of states $\textit{A}=<a_{1}=(v^{a}_{0},0),...,a_{j}=(v_{j},t^{a}_{j}),...,a_{k}=(v_{\infty},t^{a}_{k}\leq t_{max})>$ signifies the attacker’s pure strategy. Each state $a_{j}=(v_{j},t^{a}_{j})$ indicates that at time $t^{a}_{j}$, the attacker is present at node $v_{j}$. Likewise, a defender’s state $d_{r}$ is a tuple $S^{r}=(v^{r},t^{r,in},t^{r,out})$, representing the state where defender $d_{r}$ is present at the node $v^{r}$ during the interval $[t^{r,in},t^{r,out}]$.
The defender’s pure strategy consists of a set of $m$ schedules corresponding to the total number of defenders, $m$. The schedule for defender $d_{r}$ is defined as a sequence of states $S^{r}=<s^{r}_{1},...,s^{r}_{i},...,s^{r}_{k}>$, where $s^{r}_{1}=(v^{r}_{0},0,t^{r,out}_{1})$ and $s^{r}_{k}=(v^{r}_{k},t^{r,in}_{k},t_{max})$. The mixed strategy for the defender is denoted by $x=<x_{S}>$, where $x_{S}$ represents the probability with which the strategy $S$ is employed.
@@ 3 Problem description
$U_{d}\;(x,A)=\sum\limits_{S\in\textbf{S}}U_{d}(S,A)\;x_{s}$ (2)
$U_{d}\;(x,y)=\sum\limits_{A\in\textbf{A}}U_{d}(x,A)\;y_{A}$ (3)
Therefore, we need an optimization model so that the defenders can find their best optimal strategy by solving the following linear problem (see Eqs. 4-6).
@@
$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in A^{{}^{\prime}}$ (5)
$\sum\limits_{S\in\textbf{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in S^{{}^{\prime}}$ (6)
@@
First, considering that the defender commits to a strategy $x\in X$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
$BR(x)=\;argmax_{y\in Y}\;U_{a}\;(x,y)$ (7)
Second, to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
$max_{x\in S}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
Third, in a SSE, we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
$max_{x\in S,\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
SSE is popular because it is always guaranteed to exist ([21]). The flowchart of the Stackelberg game model is shown in Fig. 1.
Figure 1: Flowchart of Stackelberg game model.
@@ 4 Proposed mixed approach on multi-layer networks
This section presents the proposed methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower. In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy. We use the concept of a multi-layer network (MLN) in which, at each time-stamp, we create a copy of the entire network. We consider edge length as the time factor to create connections between these multi-layer networks. This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected. In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network. We consider a set of strategies for both players, i.e., the defenders and the attacker. For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one. For the attacker, we consider a pure strategy set.
To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies. We check the presence of each node in all defender strategies. If the node is present in multiple strategies, we sum the mixed probabilities of those strategies. This final probability is then assigned as the weight of the node. Then, we assign the weight of each node to all of its incoming edges. Finally, we apply Dijkstra’s algorithm on this network to find the optimal strategy for the attacker. For the defender, we create a multi-layer network in the same way as for the attacker. Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability. We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy. Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy. For this, we use different colors to represent each attacker strategy. The developed defender strategy includes as many different colored vertices as possible. The algorithm converges if for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$. Then the best defender strategy with the maximum probability of interdiction and total runtime are returned (see Algorithm 1).
@@ 4.1 Efficient attacker strategy design using exact approach on time expanded network
The optimal attacker strategy is generated using the exact approach for the attacker (see Algorithm 2). For each time-stamp, we create a copy of the entire network. Then, we establish connections in this multi-layer network using the edge lengths in the original graph. This algorithm describes how we assign the probability of interdiction to each node. Finally, we apply Dijkstra’s algorithm on this multi-layer network to generate the optimal strategy for the attacker.
@@
       Generate one copy of the original graph/network, labelled as $G_{i+1}$ ;
@@
Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
@@
Initial probability of interdiction $(P)$ of all nodes in all layered graphs = 0 ;
@@
                   Select node $n$ of graph $G_{t_{i}n}$ Update the minimum Prob. of Interdiction of that node $n:(P=P+P_{mix})$ where $P_{mix}$ is the mixed prob. of the current defender strategy. $t_{in}=t_{in}+1$
@@
Update the minimum Prob. of. Interdiction $(1-P)$ value to the node depending on its presence in the defender strategy:
@@
       if Prob. of interdiction of a node $n$ is $P$ then
@@
return $U_{d}$.
@@ 4.2 Efficient defender strategy design using approximation algorithm on time expanded network
If we cannot find a path that covers all colors, we can create a tree (or a set of paths) instead, which corresponds to a non-deterministic strategy for a defender.
@@ 5.2 MILP for defender: bestDo
In this context, $s^{r}_{i,v}=1$ signifies that defender $d_{r}$ reaches node $v$ at the $i^{th}$ state of $S^{r}$. Equation (23) specifies that the initial position of $d_{r}$ is $v^{r}_{0}$, and $d_{r}$ remains on a single node at each state of $S^{r}$. In Equations (24) and (25), the $0/1$ variable $\omega_{r,i,(v,u)}$ indicates whether a path exists between nodes $v$ and $u$ between the $i^{th}$ and $(i+1)^{th}$ states of $S^{r}$. Equation (26) specifies that the strategy begins at time 0 and ends at $t_{max}$, where $L_{max}$ represents the maximum allowable length of the defender’s strategy sequence, and the time the defender remains at the $i^{th}$ state is $k_{r,i}\delta$. Equation (27) guarantees that the defender travels between stops along the shortest path. With Eqs.(28)-(32), $z_{A}$ indicates whether the attacker using path $A$ is intercepted.
Specifically, $(v^{A}_{j},t^{A}_{j})$ represents the attacker’s $j^{th}$ state in $A$, while $\gamma^{A,j}_{r,i}$ indicates whether the attacker using path $A$ encounters defender $d_{r}$ at the attacker’s $j^{th}$ position while $d_{r}$ is at their $i^{th}$ stop. Additionally, $\alpha^{A,j}_{r,i}$ and $\beta^{A,j}_{r,i}$ denote whether the attacker arrives at their $j^{th}$ position after $d_{r}$ arrives or before $d_{r}$ leaves, respectively. The constant $M$ is a large number used in the formulation.
@@
To assess optimality, we compare the utility and computational time of the developed MLN-EIGS with that of MILP-EIGS.
@@ 7 Conclusion
In this paper, we present a Stackelberg game model to address security challenges in a dynamic crime scenario. We propose a novel optimal routing and resource allocation approach on a transportation network. The attacker’s optimal strategy is determined using an exact method, which, to our knowledge, is the first to apply the multi-layer network concept in solving a security game. This exact approach combines Dijkstra’s algorithm with the layered graph model. The defenders’ near-optimal strategy is generated through a novel approximation algorithm, utilizing the layered graph concept on a transportation network in a time-efficient way.
We solve the developed Stackelberg game using the adopted MILP approaches to provide optimal solutions. For benchmarking, we compare the results of our approach with those generated using MILP methods. The key contribution of this work is that our MLN-EIGS approach delivers high-quality solutions comparable to MILP-EIGS but in significantly less time.
There are some limitations, such as the exclusion of real-time traffic. Future research could focus on developing efficient algorithms to incorporate traffic data and on creating an exact approach for the defender to generate optimal strategies using the time-layered network concept.
Acknowledgements We are grateful to the members of the Multi-Agent Laboratory at Kyushu University for their insightful discussions and comments. This research is funded by a project supported by the Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science.

### v3

@@ 1 Introduction
We consider a Stackelberg game where the defender is the leader, and the attacker is the follower. The assumption is that the defenders have information only about the crime location in a large transportation network. The defenders attempt to capture the attacker before he can flee the city. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model. This mixed approach employs an exact method for generating the attacker’s strategy and an approximation method for determining the defenders’ strategies. For the defender, the developed approximation algorithm finds the efficient movement of the defender from a random initial location. For the attacker, the initial strategy is any random path from the crime location to any exit point. An exact algorithm is developed to generate the optimal path for the attacker using Dijkstra’s algorithm on the time-expanded network. The new defender strategy is generated using a greedy algorithm on the time-expanded network. The defender’s mixed strategy set is updated based on the recalculated utility. The defender then selects the best strategy to commit to. This best defender strategy is given as input to the attacker. The attacker tries to generate a new optimal strategy using the exact approach, taking the new defender strategy as input. The attacker updates the pure attacker strategy set, where all strategies are of equal probability. The algorithm checks if, for a certain number of consecutive iterations, the game utility is below a threshold value. If this convergence criterion is met, the algorithm converges, and we return the best defender strategy with the probability of interdiction and total runtime. To set a benchmark, the developed Stackelberg game is solved using the adopted MILP approaches for both the attacker and defender.
@@
Due to the complex transportation network and limited police resources, generating efficient strategies for both players is a challenging task. This paper introduces a Stackelberg game model for dynamic crime scenarios, presenting an exact attacker strategy based on a multi-layer network and a novel approximation algorithm for the defenders. The proposed MLN-EIGS approach produces high-quality solutions comparable to MILP-based methods, while significantly reducing computational time.
@@ 2 Related work
Security games play an important role in providing social security ([21], [22]). Recent literature focuses on Stackelberg games considering the escape interdiction problem to decrease the crime rate in society. For example, [4] consider search games (SEG) on directed graphs. They consider multiple defender resources and one attacker, where the attacker aims to reach one of several predefined target points from a fixed location. They develop a genetic algorithm-based heuristic approach to provide a near-optimal solution on synthetic datasets. [9] consider a leader-follower game and formulate the problem as a mathematical programming model. They use optimization software tools to solve the model by generating the optimal strategies for both leader and follower. Again, [11] consider security games on graphs and develop a polynomial-time algorithm to generate optimal strategies for players. Similarly, [10] consider security games on graphs and develop an algorithm to reduce the graph by eliminating unnecessary edges, providing a time-efficient, scalable near-optimal solution. In the same vein, [18] consider an evader-defender Stackelberg game model and develop a Monte Carlo Tree Search approach to provide efficient patrolling schemes. In addition, [23] introduce a repeated Stackelberg security game that incorporates a cooperative human behavior model to enhance patrolling strategies for wildlife protection by modeling human decision-making in repeated interactions, thereby improving defender effectiveness against adaptive adversaries in dynamic environments.
@@
3 Problem description and modeling
We consider a two-player Stackelberg game, where the sequential interaction occurs between multiple defenders $\overline{D}=\{d_{r}\mid r\in R\}$ and a single attacker $\overline{A}$. The total number of defenders is $m$, and the set of all defenders is represented by $\overline{D}$. Here, $r\in R=\{1,\dots,m\}$. Since the defenders act jointly and a complete defender strategy comprises the strategies of all individual defenders, while there is only one attacker, we refer to this as a two-player Stackelberg game.
The defenders have a finite set of actions $\acute{S}$, and the attacker has a finite set of actions represented by $\acute{A}$. The game is modeled as a general-sum game, where $x$ and $y$ are the probability distributions over $\acute{S}$ and $\acute{A}$, respectively.
The transportation network is represented as a directed graph $G=(V,E)$, where $E$ is the set of directed edges corresponding to roads, and $V$ is the set of nodes representing intersections. There is a set of predefined exit points in the considered network. $v_{\infty}$ signifies any exit node in the considered network. The game begins at time 0 and ends at time $t_{\text{max}}>0$.
The sequence of states $A=\langle a_{1}=(v^{a}_{0},0),\ldots,a_{j}=(v_{j},t^{a}_{j}),\ldots,a_{k}=(v_{\infty},t^{a}_{k}\leq t_{\text{max}})\rangle$ represents the pure strategy of the attacker. Each state $a_{j}=(v_{j},t^{a}_{j})$ indicates that at time $t^{a}_{j}$, the attacker is present at node $v_{j}$. Likewise, a defender’s state $d_{r}$ is a tuple $S^{r}=(v^{r},t^{r,in},t^{r,out})$, representing the state where defender $d_{r}$ is present at the node $v^{r}$ during the interval $[t^{r,in},t^{r,out}]$.
The defender’s pure strategy is denoted by $S$. A pure strategy for the defender consists of $m$ schedules, i.e., $S=\{S^{r}:r\in R\}$. The schedule for defender $d_{r}$ is defined as a sequence of states $S^{r}=<s^{r}_{1},...,s^{r}_{i},...,s^{r}_{k}>$, where $s^{r}_{1}=(v^{r}_{0},0,t^{r,out}_{1})$ and $s^{r}_{k}=(v^{r}_{k},t^{r,in}_{k},t_{max})$. The mixed strategy for the defender is denoted by $x=<x_{S}>$, where $x_{S}$ represents the probability with which the strategy $S$ is played.
@@ 3 Problem description and modeling
$U_{d}\;(x,A)=\sum\limits_{S\in\acute{S}}U_{d}(S,A)\;x_{s}$ (2)
$U_{d}\;(x,y)=\sum\limits_{A\in\acute{A}}U_{d}(x,A)\;y_{A}$ (3)
Here, we define an optimization model that enables the defenders to determine their optimal strategy by solving the following linear program (see Eqs. 4–6).
@@
$s.t.\qquad U^{*}\leq U_{d}\;(x,A)\qquad\forall A\in\acute{A}$ (5)
$\sum\limits_{S\in\acute{S}}x_{s}=1,x_{s}\geq 0\qquad\forall S\in\acute{S}$ (6)
@@
First, considering that the defender commits to a strategy $x\in\acute{S}$, and given such an $x$, the attacker chooses his strategy from the best-response set $BR(x)$ where
$BR(x)=\;argmax_{y\in\acute{A}}\;U_{a}\;(x,y)$ (7)
Given $x$ and $y$, the expected utility of the attacker is defined as $U_{a}(x,y)$. Second, to maximize utility, the defender chooses the strategy $x$ (the best response of the defender) to the attacker’s best response.
$max_{x\in\acute{S}}\;U_{d}\;(x,y)\qquad s.t.\;y\in BR(x)$ (8)
Third, in case of a Strong Stackelberg Equilibrium (SSE), we assume that the follower (attacker) breaks ties in favor of the leader (defender) (see Eq. 9). In that case, the optimization problem is
$max_{x\in\acute{S},\;y\in BR(x)}\;U_{d}\;(x,y)$ (9)
SSE is popular because it is always guaranteed to exist ([20]).
@@ 4 Proposed mixed approach on multi-layer networks
This section presents the proposed solution methodology for the considered escape interdiction problem which is formulated as a Stackelberg game model. A novel mixed approach on layered graph concept is developed to solve the Stackelberg model, where the defender is the leader and the attacker is the follower. In this Stackelberg game, first, the defenders commit to a strategy, and then the attacker generates the best response to the given defender strategy. We use the concept of a multi-layer network (MLN) in which, at each time-stamp, we create a copy of the entire network. We consider edge length as the time factor to create connections between these multi-layer networks. This means that depending on the edge length, we choose the layers from which the start and end nodes of that particular edge are selected. In this way, for all edges in the original transportation network, we create corresponding edges in the multi-layer network. We consider a set of strategies for both players, i.e., the defenders and the attacker. For the defender, we use a mixed strategy set where each strategy is assigned a mixed probability, with the sum of these probabilities equaling one. For the attacker, we consider a pure strategy set.
To generate the optimal attacker strategy, we first assign weights to the nodes depending on the mixed probabilities and the nodes’ presence in the defender strategies. We check the presence of each node in all defender strategies. If the node is present in multiple defender strategies, we sum the mixed probabilities of those strategies. This resulting probability $P$ is then assigned as the weight of the node. Subsequently, a weight of $P$ is assigned to all of its incoming edges, enabling Dijkstra’s algorithm to identify the attacker’s optimal path with the minimum probability of interdiction.
CoreLP computes the mixed strategies of the defender and attacker, $x$ and $y$, by solving the linear program defined in Eqs. (4)–(6), using the complete strategy spaces of the defender and attacker, denoted by $\acute{S}$ and $\acute{A}$, respectively, as input. Here, $ExactAO$ denotes the exact approach developed for the attacker, while $ApproxDO$ denotes the approximation algorithm developed for the defender.
For the defender, we create a multi-layer network in the same way as for the attacker. Weights are assigned to all nodes based on the attacker strategies, with each attacker strategy given equal probability. We develop a novel approximation algorithm for the defender to generate a near-optimal defender strategy. Thus, for the defender, the problem can be considered as finding a near-optimal path with the aim of covering at least one node from each attacker strategy. For this, we use different colors to represent each attacker strategy. The developed defender strategy includes as many different colored vertices as possible. The algorithm converges if for $K$ consecutive iterations, the game utility is below a threshold value named $\epsilon$. Then the best defender strategy with the maximum probability of interdiction and total runtime are returned (see Algorithm 1).
@@
The flowchart of the Stackelberg game model is shown in Fig. 1.
Figure 1: Flowchart of Stackelberg game model.
@@ 4.1 Efficient attacker strategy design using exact approach on time expanded network
The optimal attacker strategy is computed using an exact approach (see Algorithm 2). For each time-stamp, a separate copy of the entire network is created, forming a multi-layer structure. Connections between these layers are established based on the edge lengths in the original graph. Initially, all edge weights in the multi-layer network are set to zero. The algorithm outlines the process of assigning interdiction probabilities to each node and the corresponding weights to its adjacent edges. Finally, Dijkstra’s algorithm is applied to this multi-layer network to determine the attacker’s optimal strategy.
@@
       Generate one copy of the original graph/network, labeled as $G_{i+1}$ ;
@@
Update the maximum probability of interdiction for each node:
@@
Initial probability of interdiction $(P)$ of all nodes in the layered graph = 0 ;
Initial weight of all edges in the layered graph = 0 ;
@@
                   Select node $n$ of graph $G_{t_{in}}$;
                   Update the maximum probability of Interdiction of that node $n:(P=P+P_{mix})$ where $P_{mix}$ is the mixed prob. of the current defender strategy.;
                   $t_{in}=t_{in}+1$;
@@
Update the weights of all incoming edges to the node with interdiction probability $P$ as $P$:
@@
       if Probability of interdiction of a node $n$ is $P$ then
@@
return The optimal attacker path with the minimum probability of interdiction.
@@ 5.2 MILP for defender: bestDo
In this context, $s^{r}_{i,v}=1$ signifies that defender $d_{r}$ reaches node $v$ at the $i^{th}$ state of $S^{r}$. Equation (23) specifies that the initial position of $d_{r}$ is $v^{r}_{0}$, and $d_{r}$ remains on a single node at each state of $S^{r}$. In Equations (24) and (25), the $0/1$ variable $\omega_{r,i,(v,u)}$ indicates whether a path exists between nodes $v$ and $u$ between the $i^{th}$ and $(i+1)^{th}$ states of $S^{r}$. Equation (26) specifies that the strategy begins at time 0 and ends at $t_{max}$, where $L_{max}$ represents the maximum allowable length of the defender’s strategy sequence, and the time the defender remains at the $i^{th}$ state is $k_{r,i}\delta$. Equation (27) guarantees that the defender travels between stops along the shortest path. With Eqs.(28)-(32), $z_{A}$ indicates whether the attacker using path $A$ is intercepted. Specifically, $(v^{A}_{j},t^{A}_{j})$ represents the attacker’s $j^{th}$ state in $A$, while $\gamma^{A,j}_{r,i}$ indicates whether the attacker using path $A$ encounters defender $d_{r}$ at the attacker’s $j^{th}$ position while $d_{r}$ is at their $i^{th}$ stop. Additionally, $\alpha^{A,j}_{r,i}$ and $\beta^{A,j}_{r,i}$ denote whether the attacker arrives at their $j^{th}$ position after $d_{r}$ arrives or before $d_{r}$ leaves, respectively. The constant $M$ is a large number used in the formulation.
@@
To evaluate the efficacy of our proposed approach, MLN-EIGS, we compare its utility and computational time with those of the adopted method, MILP-EIGS.
@@ 7 Conclusion
We address a Stackelberg game formulation of the escape interdiction game (EIG) using the multi-layer network concept, focusing on the development of effective solution methodologies. An exact approach is proposed for the attacker and a near-optimal approach for the defenders. To the best of our knowledge, this is the first implementation of the escape interdiction problem using a multi-layer network framework. Our paper presents a proof establishing the NP-completeness of the ’color covering’ problem. Using grid datasets, we demonstrate that our method significantly outperforms the MILP approach in both computational time and solution quality, with minimal performance trade-offs.
Some limitations remain, such as the exclusion of real-time traffic data. Future research could aim to develop efficient algorithms that incorporate traffic dynamics and to design an exact approach for the defender to generate optimal strategies using the time-layered network model.
Acknowledgments We are grateful to the members of the Multi-Agent Laboratory at Kyushu University for their insightful discussions and comments. This research is funded by a project supported by the Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science.